Find the area of the region bounded by the curves X= Y³ and Y= X²

1
Expert's answer
2022-05-26T17:19:25-0400
"y=\sqrt[3]{x}, y=x^2"

"\sqrt[3]{x} =x^2"

"x=x^6"

"x(x^5-1)=0"

"x_1=0, x_2=1"

"A=\displaystyle\int_{0}^{1}(\sqrt[3]{x} -x^2)dx""=\big[\dfrac{3x^{4/3}}{4}-\dfrac{x^3}{3}\big]\begin{matrix}
 1 \\
 0
\end{matrix}""=\dfrac{3}{4}-\dfrac{1}{3}=\dfrac{5}{12}({units}^2)"